# Structured Sparse Transition Matrices to Enable State Tracking in State-Space Models

### v2

@@ 4.4 Long-Range Arena
The long-range arena [Tay et al. 2021] dataset (LRA) covers mathematical expressions, freeform text, and images, all represented in sequences up to length 16k. We evaluate our model on a subset of LRA, with sequences up to length 4k. The baseline results are taken from Soydan et al. [2024]. We consistently use 4 layers with embedding dimension 128 and state dimension 128 for all tasks except Retrieval, which used state dimension 64. For all tasks, the transition matrix dictionary size is set to $K=6$. Time-invariant SSMs perform significantly better on average than the shown collection of time-variant ones. Among the time-variant SSMs (Mamba, S7, and PD-SSM), ours performs best on average on this set of tasks. Together with the time-series results, this serves as evidence that the PD parametrization can be effective on realistic long-sequence tasks. Time-Invariant Time-Variant Dataset S4 LRU     Mamba S7 PD-SSM ListOps 59.6 60.2     38.0 63.8 61.0 Text 86.8 89.4     83.0 87.2 88.1 Image 88.6 89.0     69.8 61.1 70.4 Retrieval 90.9 89.9     72.1 91.8 90.0 Pathfinder 94.2 95.1     69.3 65.5 62.6 Average 84.0 84.7     66.4 73.9 74.4 Table 5: LRA results, average over 3 seeds.
@@ 4.5 State Tracking in Natural Language
Figure 5: Cayley diagram of the $A_{5}$ group [Carter 2009]. We encode state transitions into sets of English sentences.
@@
oding a transportation network where an agent can either move along a ring of motorized transportation (via bus, plane, train, car) or move (i.e., walk, run, jog, hike) between such rings. Red transitions correspond to movement using motorized transport, whereas blue transitions are taken without motorized transport, see Figure 5. Figure 6 shows the performance of diagonal SSMs as well as PD-SSM for state tracking in natural language. To benefit from large-scale pretraining and to demonstrate the modularity of our approach, we freeze a pretrained Qwen 2.5-1.5B model and replace its final layer with a single trainable SSM layer. This analysis centers around the applicability of the sparse parametrization in a larger-scale setting, and also serves as a test of the utility of the $PD$ parametrization as opposed to purely diagonal models. Concretely, we vary the matrix structure such that it is either real-valued diagonal ($A(u_{t})=|D(u_{t})|$), complex-valued diagonal ($A(u_{t})=D(u_{t})$), or $A(u_{t})=P(u_{t})D(u_{t})$. We train the model for 100,000 steps with batches of size 256. Each batch consists of sequences of English sentences, with the number of sentences randomly sampled up to 25, with each sentence triggering a state transition in the underlying automaton. The automata are Parity and $A_{5}$ with two generators as in Figure 5. In Figure 6 we report the best validation performance out of 3 random seeds of each model with an equivalent hyperparameter grid search. Even though the complex diagonal model can represent parity [Sarrof et al. 2024], it does not learn to do so in practice, while PD-SSM converges to the generalizing solution within 10,000 steps.
@@
Figure 6: Applying one layer of our PD-SSM at the output of a frozen Qwen-2.5 1.5B allows the model to learn to follow the state transitions of FSAs of various complexities. The vertical dashed line indicates the maximum training sequence length.
@@ 5 Conclusion
We have introduced an expressive yet efficient parametrization of transition matrices for selective state-space models manifesting in the PD-SSM architecture. Theoretically, PD-SSMs can emulate arbitrary finite-state automata while reaching the lower bound on the state size required to do so in the worst case. They are significantly more efficient than the similarly performant unstructured SD-SSM, and reach a new state-of-the-art in synthetic state-tracking tasks. We verifed its effectiveness on a novel state-tracking in natural language benchmark. Among the investigated time-variant SSMs, PD-SSM on average performs best on tasks requiring the processing of long sequences (up to 4k), and exhibits high performance on time-series classification with long sequences.
@@
Although we significantly improve upon the scalability of SD-SSMs while retaining its favorable expressivity, PD-SSMs still incur overhead when generating the $P(u_{t})$ matrices. Our future work will investigate how PD-SSM can benefit from more efficient one-hot column matrix generation strategies, and how it can be utilized in large-scale pretraining.
@@ 6 Broader Impact
The paper introduces a novel neural network architecture with increased expressiveness and efficiency. More expressive and efficient models can potentially have societal impacts in the future, but we do not see any immediate and direct negative impact of this work.
@@ Acknowledgement
This work is also supported by the Swiss National Science foundation (SNF), grant 10002666.
@@ Appendix A Background
The goal of this section is to provide knowledge of abstract algebra and circuit complexity theory at a level sufficient to understand the motivation behind our paper. While the presentation is fully precise, we do omit certain details for clarity, and we often focus on providing an intuitive understanding by considering visual and textual explanations and examples for important concepts. This approach is highly inspired by Carter [2009]. As the core of our work is the design and analysis of a novel neural network architecture, one does not need to have a deep understanding of these topics to understand our paper. As such, the entire Appendix A can be skipped without significantly affecting the understanding of our PD method. For a deeper treatment of the topics, one may consult the following resources: [Straubing 1994, Carter 2009, Arora and Barak 2009].
@@ A.1 Algebra and Finite-State Automata
Figure S1: Left: A visualization of the $A_{5}$ automaton as we used it in our experiments [Carter 2009]. Center: The two transitions of $A_{5}$ equivalently represented in matrix form. $\mathcal{T}(\mathcal{A})$ is isomorphic to the set of all possible unique iterated products of the two matrices, of which there are a total of 60. Right: A simplified diagonal SSM that omits the $B(u_{t})u_{t}$ term. As the transition matrices are restricted to be diagonal, and since the two transition matrices represented in the center are not simultaneously diagonalizable, there seems to exist a discrepancy between $\mathcal{T}(\mathcal{A})$ and the SSM.
The reasoning in the figure above is only valid for a single SSM layer with no $B(u_{t})u_{t}$ term. In order to reason about several layers of the full model, recent work Merrill et al. [2024], Sarrof et al. [2024] have utilized the frameworks of Abstract Algebra (A.2) and Circuit Complexity Theory (A.3).
@@ A.2 Algebraic Groups and their Properties
Figure S2: Cayley diagrams of the group $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ and the dihedral group $D_{4}$ [Carter 2009]. $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ is a commutative group, while $D_{4}$ is not. Both are solvable, with solvability of $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ shown in Example 3.
@@
Another central property of groups is solvability. We will first present it informally and will then precisely define it. Loosely speaking, a group is solvable if it can be constructed from co- [Uncaptioned image] Figure S3: Composition of $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$. mmutative groups. The details of the construction are particularly easy to understand on the example of the $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ group, shown in Figure S3. The $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ group is solvable, as it corresponds to the direct product of two commutative groups. While the $D_{4}$ group shown previously is not commutative, it is solvable as it can be constructed from commutative groups in a similar manner as $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$. However, $A_{5}$ cannot be constructed using commutative groups Carter [2009]. To speak about solvability in more definite terms, we need to define several concepts. The re- mainder of this subsection defines and explains solvability in concrete terms. It can be freely skipped without any effect on reading the rest of the appendix. Firstly, we will define subgroups, which are in essence closed subsets of a group.
@@
Of central interest are normal subgroups, which is a particularly well-behaved category of subgroup. We will normal subgroups them to partition a group into a sequence of subgroups. The solvability of a group can be determined by examining the structure of a sequence of normal subgroups.
@@
Let us again consider the even and odd integers. We have seen that this is the full set of cosets of $\mathbb{Z^{+}}$. If we abstract away the individual numbers and only consider the concepts of even and odd, we can see that the two concepts form a group structure.
@@
We are now equipped with a basic understanding of what a solvable group is. To better understand the implications of this concept for state-tracking with SSMs, we need to introduce a different, yet perhaps surprisingly related framework [Straubing 1994], that of Boolean circuit complexity.
@@ A.3 Circuit Complexity Theory
The theory of circuit complexity aims to provide an abstract understanding of highly parallel algorithms. One might imagine that there exist problems for which it is difficult, if not impossible, to design one efficient algorithm which operates on arbitrary-length inputs, yet in which all input lengths admit a bespoke, highly efficient and parallel algorithm [^2] As per [Arora and Barak 2009], introductory remark to Chapter 6, Karp and Lipton propose a variant of this idea in 1982.. Such considerations are especially important in cryptography, as an efficient algorithm for the factorization of n-bit integers would be a significant development for any large $n$. Boolean circuit complexity is the study of such parallel algorithms, or rather circuit families. Let us first define a Boolean circuit.
@@
$TC^{0}$ consists of shallow (constant-depth) but arbitrary fan-in Boolean circuits of standard elements (AND, OR, NOT) augmented with a majority gate. Emulating solvable automata is proven to be within $TC^{0}$ [Sarrof et al. 2024].
@@
$NC^{1}$ allows for logarithmic depth, but restricts the fan-in of the gates to 2. Emulating solvable automata is proven to be an $NC^{1}$-complete problem Barrington [1989], meaning that all $NC^{1}$ problems can be reduced to it.
Many relations between circuit complexity classes remain unknown. It is, for example, not known whether $TC^{0}\neq NC^{1}$. It is in fact not even known whether $TC^{0}$ is equal to $P$, the class of problems decidable in polynomial time [Arora and Barak 2009]. As a consequence, certain restrictions have been placed on the generation of Boolean circuits in order to allow for a more realistic and fine-grained study of their properties. The central restriction is that of uniform generation, or uniformity.
@@ A.4 SSM Expressivity Results
We finally arrive at the results that this section was building up to, bounds on the expressivity of SSMs within the framework of algebraic groups and Boolean circuits. We do not prove any new theorems here nor restate the proofs, opting instead to refer to work which proves the claims. Let us start by restating a classic result in the terminology that we introduced in our paper:
Theorem 1 (Barrington [1989], Theorem 5).
The word problem of a group is defined as follows Merrill et al. [2024]:
@@
The separation of word problems over solvable and non-solvable groups rests on the widely-held conjecture that $TC^{0}\neq NC^{1}$ Arora and Barak [2009], combined with the result that word problems of solvable groups belong to $TC^{0}$:
Theorem 2 (Mix Barrington et al. [1992], Theorem 8).
@@
A recent work analyses the expressivity of diagonal SSMs by demonstrating that the individual computations in the SSM can be emulated by $\text{logspace-uniform }TC^{0}$ circuits, under the assumption that the numerical representation capacity is logarithmic in the sequence length. Concretely, it is $O(\log n)$.
Theorem 3 (Merrill et al. [2024], Theorem 4.4).
@@
This is the main upper bound we refer to in the paper, and is what we mean when we say that non-solvable automata cannot be emulated by diagonal SSMs, which is implied by combining the strict inclusion logspace-uniform $TC0\displaystyle\text${logspace-uniform $\subset TC^{0}$ with the conjectured $TC^{0}\cap NC^{1}\neq\emptyset$, and utilizing Theorem 3. This bound rests on an unproven conjecture, but it is supported by significant experimental evidence Merrill et al. [2024], Cirone et al. [2024], Grazzi et al. [2024], Sarrof et al. [2024], Terzić et al. [2025] [^3] Using tools from rough path theory[Cirone et al. 2024] prove the restrictiveness of diagonal time-variant SSMs as compared to dense ones without relying on unproven circuit complexity conjectures..
@@ B.1 Full Architecture
This work mainly describes a new method for generating the SSM transition matrices. The method is itself embedded into a larger architecture, which mostly follows standard practice in neural network design as pioneered by the Transformer [Vaswani et al. 2017]. That is, we embed the $PD$ parametrization into neural network layers which consist of nonlinear feed-forward networks, vector normalization, and residual connections, calling the entire model PD-SSM. The connection pattern of PD-SSM closely follows that of pre- or post-norm Transformers, and can be seen in [Xiong et al. 2020]. We show two sketches of the full PD-SSM model in Figure S9.
@@
The readout from the state, $\psi(x_{t})$, is different from standard practice, which typically implements it as $\psi(x_{t})=Re\{x_{t}\}$ [Orvieto et al. 2023, Gu et al. 2022b, Gu and Dao 2023, Smith et al. 2023]. Instead of taking the real part of $x_{t}$, we apply a transformation on the concatenation of the real and imaginary part of $x_{t}$, denoted as $Re\{x_{t}\}||Im\{x_{t}\}$. Depending on the benchmark, the readout is either a GELU-activated nonlinear network:
@@
Or, if we refer to it as a linear readout [^4] It is typical terminology to refer to trainable affine transformations as linear neural network layers., we concretely mean the affine transformation:
@@
Figure S9: Full PD-SSM models we used in our state-tracking and LRA experiments. In the time-series classification experiments, we used a model equivalent to the one shown in subfigure 9(b), with the exception that $\tilde{x}_{t}=Re\{x_{t}\}\in\mathbb{R}^{N}$.

### v3

@@ 4.4 Long-Range Arena
The long-range arena [tay2020long] dataset (LRA) covers mathematical expressions, freeform text, and images, all represented in sequences up to length 16k. We evaluate our model on a subset of LRA, with sequences up to length 4k. The baseline results are taken from soydan_2024_simplified. We consistently use 4 layers with embedding dimension 128 and state dimension 128 for all tasks except Retrieval, which used state dimension 64. For all tasks, the transition matrix dictionary size is set to $K=6$. Time-invariant SSMs perform significantly better on average than the shown collection of time-variant ones. Among the time-variant SSMs (Mamba, S7, and PD-SSM), ours performs best on average on this set of tasks. Together with the time-series results, this serves as evidence that the PD parametrization can be effective on realistic long-sequence tasks. Time-Invariant Time-Variant Dataset S4 LRU     Mamba S7 PD-SSM ListOps 59.6 60.2     38.0 63.8 61.0 Text 86.8 89.4     83.0 87.2 88.1 Image 88.6 89.0     69.8 61.1 70.4 Retrieval 90.9 89.9     72.1 91.8 90.0 Pathfinder 94.2 95.1     69.3 65.5 62.6 Average 84.0 84.7     66.4 73.9 74.4 Table 5: LRA results, average over 3 seeds.
@@ 4.5 State Tracking in Natural Language
Figure 5: Cayley diagram of the $A_{5}$ group [carter_visual_2009]. We encode state transitions into sets of English sentences.
@@
oding a transportation network where an agent can either move along a ring of motorized transportation (via bus, plane, train, car) or move (i.e., walk, run, jog, hike) between such rings. Red transitions correspond to movement using motorized transport, whereas blue transitions are taken without motorized transport, see Figure 5. Figure 6 shows the performance of diagonal SSMs as well as PD-SSM for state tracking in natural language. To benefit from large-scale pretraining and to demonstrate the modularity of our approach, we freeze a pretrained Qwen 2.5-1.5B model and replace its final layer with a single trainable SSM layer. This analysis centers around the applicability of the sparse parametrization in a larger-scale setting, and also serves as a test of the utility of the $PD$ parametrization as opposed to purely diagonal models. Concretely, we vary the matrix structure such that it is either real-valued diagonal ($A(u_{t})=|D(u_{t})|$), complex-valued diagonal ($A(u_{t})=D(u_{t})$), or structured sparse ($A(u_{t})=P(u_{t})D(u_{t})$). We train the model for 100,000 steps with batches of size 256. Each batch consists of sequences of English sentences, with the number of sentences randomly sampled up to 25, each sentence triggering a state transition in the underlying automaton. The automata are Parity and $A_{5}$ with two generators as in Figure 5. In Figure 6 we report the best validation performance out of 3 random seeds of each model with an equivalent hyperparameter grid search. Even though complex diagonal SSMs can represent parity [sarrof2024expressivecapacitystatespace], the models did not learn to do so in the experiments. On this task, PD-SSM converges within 10,000 steps.
@@
Figure 6: Applying one layer of our PD-SSM at the output of a frozen Qwen-2.5 1.5B allows the model to learn to follow the state transitions of automata when the inputs are redundantly encoded in English sentences. The vertical dashed line indicates the maximum training sequence length.
@@ 5 Conclusion
We introduced an expressive and efficient structured sparse parametrization of transition matrices for time-varying SSMs manifesting in the PD-SSM architecture. Theoretically, PD-SSM can emulate arbitrary finite-state automata while reaching the lower bound on the network depth and state size required to do so in the worst case. It is significantly more efficient than unstructured SSMs and reaches a new state-of-the-art in synthetic state-tracking tasks. Among the investigated time-varying SSMs, PD-SSM on average performs best on tasks from the LRA benchmark requiring the processing of long sequences (up to 4k), and additionally exhibits high performance on multivariate time-series classification with long sequences (up to 17k), demonstrating performance within standard error of the state-of-the-art SSM. Finally, we verified its effectiveness in a hybrid Attention-SSM architecture on a novel state-tracking in natural language task, showing that it can track non-solvable automaton states even when the inputs are redundantly encoded in variable-length English sentences.
@@
Although we significantly improve upon the scalability of SD-SSM while retaining its favorable expressivity, PD-SSM still incurs overhead when generating the $P(u_{t})$ matrices. Our future work will investigate how our method can benefit from more efficient one-hot column matrix generation strategies, and how it can be utilized in large-scale pretraining.
@@ 6 Broader Impact
The paper introduces a novel neural network architecture with increased expressivity while retaining efficiency. More expressive and efficient models can potentially have unpredictable societal impacts in the future, but we do not foresee any immediate and direct negative impact of this work.
@@ Acknowledgement
This work is supported by the Swiss National Science foundation (SNF), grant 10002666.
@@ Appendix A Background
The goal of this section is to provide knowledge of abstract algebra and circuit complexity theory at a level sufficient to understand the motivation behind our paper. We omit certain details for reasons of clarity, and we often focus on providing an intuitive understanding by considering visual and textual explanations and examples for important concepts. This approach is highly inspired by carter_visual_2009. As the core of our work is the design and analysis of a novel neural network architecture, one does not need to have a deep understanding of these topics to understand our paper. As such, the entire Appendix A can be skipped without significantly affecting the understanding of our PD method. For a deeper treatment of the topics, one may consult the following resources: [straubing_book, carter_visual_2009, arora_barak_2009].
@@ A.1 Algebra and Finite-State Automata
Figure S1: Left: A visualization of the $A_{5}$ automaton as we used it in our experiments [carter_visual_2009]. Center: The two transitions of $A_{5}$ equivalently represented in matrix form. $\mathcal{T}(\mathcal{A})$ is isomorphic to the set of all possible unique iterated products of the two matrices, of which there are a total of 60. Right: A simplified diagonal SSM that omits the $B(u_{t})u_{t}$ term. As the transition matrices are restricted to be diagonal, and since the two transition matrices represented in the center are not simultaneously diagonalizable, there seems to exist a discrepancy between $\mathcal{T}(\mathcal{A})$ and the SSM.
The reasoning in the figure above is only valid for a single SSM layer with no $B(u_{t})u_{t}$ term. In order to reason about several layers of the full model, recent work merrill_illusion_2024, sarrof2024expressivecapacitystatespace have utilized the frameworks of Abstract Algebra (A.2) and Circuit Complexity Theory (A.3).
@@ A.2 Algebraic Groups and their Properties
Figure S2: Cayley diagrams of the group $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ and the dihedral group $D_{4}$ [carter_visual_2009]. $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ is a commutative group, while $D_{4}$ is not. Both are solvable, with solvability of $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ shown in Example 3.
@@
Another central property of groups is solvability. We will first present it informally and will then precisely define it. Loosely speaking, a group is solvable if it can be constructed from co- [Uncaptioned image] Figure S3: Composition of $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$. mmutative groups. The details of the construction are particularly easy to understand on the example of the $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ group, shown in Figure S3. The $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ group is solvable, as it corresponds to the direct product of two commutative groups. While the $D_{4}$ group shown previously is not commutative, it is solvable as it can be constructed from commutative groups in a similar manner as $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$. However, $A_{5}$ cannot be constructed using commutative groups carter_visual_2009. To speak about solvability in more definite terms, we need to define several concepts. The re- mainder of this subsection defines and explains solvability in concrete terms. Firstly, we will define subgroups, which are in essence closed subsets of a group.
@@
Of central interest are normal subgroups, which is a particularly well-behaved family of subgroup. We will use normal subgroups to partition a group into a sequence of subgroups. The solvability of a group can be determined by examining the structure of a sequence of normal subgroups.
@@
Let us again consider the even and odd integers. We have seen that this is the full set of cosets of $\mathbb{Z^{+}}$. If we abstract away the individual numbers and only consider the concepts of even and odd, we can see that the two concepts form a group structure, with the natural operation defined as per the sketch below.
@@
We are now equipped with a basic understanding of what a solvable group is. To better understand the implications of this concept for state-tracking with SSMs, we need to introduce a different, yet perhaps surprisingly related framework [straubing_book], that of Boolean circuit complexity.
@@ A.3 Circuit Complexity Theory
The theory of circuit complexity aims to provide an abstract understanding of highly parallel algorithms. One might imagine that there exist problems for which it is difficult, if not impossible, to design one efficient algorithm which operates on arbitrary-length inputs, yet in which all input lengths admit a bespoke, highly efficient and parallel algorithm [^3] As per [arora_barak_2009], introductory remark to Chapter 6, Karp and Lipton propose a variant of this idea in 1982.. Such considerations are especially important in cryptography, as an efficient algorithm for the factorization of n-bit integers would be a significant development for any large $n$. Boolean circuit complexity is the study of such parallel algorithms, or rather circuit families. Let us first define a Boolean circuit.
@@
$TC^{0}$ consists of shallow (constant-depth) but arbitrary fan-in Boolean circuits of standard elements (AND, OR, NOT) augmented with a majority gate. Emulating solvable automata is proven to be within $TC^{0}$ [sarrof2024expressivecapacitystatespace].
@@
$NC^{1}$ allows for logarithmic depth, but restricts the fan-in of the gates to 2. Emulating solvable automata is proven to be an $NC^{1}$-complete problem barrington_1989, meaning that all $NC^{1}$ problems can be reduced to it.
Many relations between circuit complexity classes remain unknown. It is, for example, not known whether $TC^{0}\neq NC^{1}$. It is in fact not even known whether $TC^{0}$ is equal to $P$, the class of problems decidable in polynomial time [arora_barak_2009]. As a consequence, certain restrictions have been placed on the generation of Boolean circuits in order to allow for a more realistic and fine-grained study of their properties. The central restriction is that of uniform generation, or uniformity.
@@ A.4 SSM Expressivity Results
We finally arrive at the results that this section was building up to, bounds on the expressivity of SSMs within the framework of algebraic groups and Boolean circuits. We do not prove any new theorems here nor restate the proofs, opting instead to refer to work which proves the claims. We will use the notion group world problem, defined as follows merrill_illusion_2024:
@@
A classic result then states:
Theorem 1 (barrington_1989, Theorem 5).
The separation of word problems over solvable and non-solvable groups rests on the widely-held conjecture that $TC^{0}\neq NC^{1}$ arora_barak_2009, combined with the result that word problems of solvable groups belong to $TC^{0}$:
Theorem 2 (barrington_nc1_1992, Theorem 8).
@@
A recent work analyses the expressivity of diagonal SSMs by demonstrating that the individual computations in the SSM can be emulated by $\text{logspace-uniform }TC^{0}$ circuits, under the assumption that the numerical representation capacity is logarithmic in the sequence length.
Theorem 3 (merrill_illusion_2024, Theorem 4.4).
@@
This is the main upper bound we refer to in the paper, and is what we mean when we say that non-solvable automata cannot be emulated by diagonal SSMs, which is implied by combining the strict inclusion logspace-uniform $TC0\displaystyle\text${logspace-uniform $\subset TC^{0}$ with the conjectured $TC^{0}\cap NC^{1}\neq\emptyset$, and utilizing Theorem 3. This bound rests on an unproven conjecture, but it is supported by significant experimental evidence merrill_illusion_2024, cirone_theoretical_2024, grazzi2024unlocking, sarrof2024expressivecapacitystatespace, terzic2025sdssm [^4] Using tools from rough path theory[cirone_theoretical_2024] prove the restrictiveness of diagonal time-variant SSMs as compared to dense ones without relying on unproven circuit complexity conjectures..
@@ B.1 Full Architecture
This work mainly describes a new method for generating the SSM transition matrices. The method is itself embedded into a larger architecture, which mostly follows standard practice in neural network design as pioneered by the Transformer [vaswani_attention_2017]. That is, we embed the $PD$ parametrization into neural network layers which consist of nonlinear feed-forward networks, vector normalization, and residual connections, calling the entire model PD-SSM. The connection pattern of PD-SSM closely follows that of pre- or post-norm Transformers, and can be seen in [pre_post_norm]. We show two sketches of the full PD-SSM model in Figure S9.
@@
The readout from the state, $\psi(x_{t})$, is different from standard practice, which typically implements it as $\psi(x_{t})=Re\{x_{t}\}$ [orvieto_resurrecting_2023, gu_efficiently_2022, gu_mamba_2023, smith_simplified_2023]. Instead of taking the real part of $x_{t}$, we apply a transformation on the concatenation of the real and imaginary part of $x_{t}$, denoted as $Re\{x_{t}\}||Im\{x_{t}\}$. Depending on the benchmark, the readout is either a GELU-activated nonlinear network:
@@
Or, if we refer to it as a linear readout, we concretely mean the transformation:
@@
Figure S9: Full PD-SSM models we used in our state-tracking and LRA experiments. In the time-series classification experiments, we used a model equivalent to the one shown in subfigure 9(a), with the exception that $\tilde{x}_{t}=Re\{x_{t}\}\in\mathbb{R}^{N}$.